Hi, I am having trouble on this question...
Write an equation of the line containing the given point and parallel to the given line. Express in y=mx+b (-4,8); 4x=3y+2
Thanks for any help. I appreciate it!!

Answer by nerdybill(7384)

Write an equation of the line containing the given point and parallel to the given line. Express in y=mx+b (-4,8); 4x=3y+2
.
First, we determine the slope of:
4x=3y+2
Do this by rearranging the equation so that it is in "slope-intercept" form:
y = mx + b
where
m is slope
b is y-intercept
.
4x=3y+2
4x-2 = 3y
(4/3)x - 2/3 = y
y = (4/3)x - 2/3
The slope then is 4/3.
.
If a two lines are parallel, their slopes MUST be the same.
Therefore our NEW line:
m = 4/3
AND, crosses;
(-4,8)
.
Plug the above into the "point-slope" form:
y - y1 = m(x-x1)
y - 8 = (4/3)(x-(-4))
y - 8 = (4/3)(x+4)
y - 8 = (4/3)x + 16/3
y = (4/3)x + 16/3 + 8
y = (4/3)x + 16/3 + 24/3
y = (4/3)x + 40/3 (this is what they're looking for)